# Copresheaf Topological Neural Networks: A Generalized Deep Learning Framework

### v2

@@ 4 Architectures Derived from the Copresheaf Framework
Having established the abstract copresheaf-based framework on a combinatorial complex $\mathcal{X}$, we now present several concrete instantiations. Copresheaf transformers (CTs) extend the standard attention mechanism by dynamically learning linear maps $\rho_{y\to x}:\mathcal{F}(y)\to\mathcal{F}(x)$ encoding directional, anisotropic relationships between tokens, i.e., cells, in $\mathcal{X}$. Integrating these maps into attention enables CTs to capture rich, structured interactions. Copresheaf graph neural networks (CGNNs) generalize message passing GNNs by incorporating copresheaf linear maps to model relational structures. Copresheaf Convolutional Networks define convolution-like operations on combinatorial complexes, modeled as Euclidean grid, using these linear maps. We present the CT construction next and leave the exact formulations of copresheaf networks, CGNNs, and copresheaf convolution layers to the appendix.
@@
Having established the abstract framework of CTNNs on a combinatorial complex $\mathcal{X}$, we now introduce a concrete instantiation: the Copresheaf Transformer (CT) layer. This layer extends the standard attention mechanism by dynamically learning linear maps $\rho_{y\to x}:\mathbb{R}^{d_{y}}\to\mathbb{R}^{d_{x}}$ that encode both the combinatorial structure and directional, anisotropic relationships within the complex. By integrating these maps into the attention computation, the CT layer captures rich, structured interactions across $\mathcal{X}$. At layer $\ell$, each cell $x\in\mathcal{X}$ is associated with a feature $\mathbf{h}_{x}^{(\ell)}\in\mathbb{R}^{d_{x}}$, where $d_{x}$ denotes the feature dimension of cell $x$.
Proposition 4 (Copresheaf self-attention).
For a fixed rank $k$ and neighborhood $\mathcal{N}_{k}$ (e.g., adjacency between $k$-cells), let $W_{q},W_{k}\in\mathbb{R}^{p\times d},\quad W_{v}\in\mathbb{R}^{d\times d}$ denote learnable projection matrices, where $p$ is the dimension of the query and key spaces, and $d$ is the feature dimension (assumed uniform across cells for simplicity). For each $k$-cell $x\in\mathcal{X}^{k}$, copresheaf self-attention defines the message aggregation and feature update as $\mathbf{h}_{x}^{(\ell+1)}=\beta\bigl{(}\mathbf{h}_{x}^{(\ell)},m_{x}\bigr{)}$, where $m_{x}=\sum_{y\in\mathcal{N}_{k}(x)}a_{xy}\rho_{y\to x}(v_{y})$ and
@@
where $q_{x}=W_{q}h_{x}^{(\ell)},k_{x}=W_{k}h_{x}^{(\ell)},~{}\mathrm{and}~{}v_{x}=W_{v}h_{x}^{(\ell)}$. Here, the the softmax normalizes over all neighbors $y^{\prime}\in\mathcal{N}_{k}(x)$ and $\rho_{y\to x}:\mathcal{F}(y)\to\mathcal{F}(x)$ is the learned copresheaf map. The update function $\beta$ is chosen to be a neural network.
Similarly, we define copresheaf cross-attention among $s$ and $t$ rank cells in $\mathcal{X}$ detailed in our appendix, as well as a general algorithm for a corpresheaf transformer layer.
@@ 5 Experimental Evaluation
We conduct synthetic and real experiments in numerous settings to support the generality of our framework. These include learning physical dynamics, graph classification in homophillic and heterophilic cases and classifying higher-order complexes.
@@ 5.1 Evaluations on Physics Datasets
To verify the validity of our networks in toy setups of different characters, we generate a series of synthetic datasets. These include:
@@
    Advection. Similar to heat, we generate 600 realisations of $u_{t}+c\,u_{x}=0$ with $c=1$; the solution is a pure phase shift of $u_{0}$, sampled on $N=130$ points. Pairs are min–max scaled to $[0,1]$ and split $500{:}100$ for train/test.
@@
    Unsteady stokes. Let $\mathbf{u}(x,y,t)=\bigl{(}u(x,y,t),\,v(x,y,t)\bigr{)}\in\mathbb{R}^{2}$ denote the incompressible velocity field, $p(x,y,t)\in\mathbb{R}$ the kinematic pressure, and $\nu>0$ the kinematic viscosity. Throughout, $\partial_{t}$ is the time derivative, $\nabla=(\partial_{x},\partial_{y})$ the spatial gradient, and $\Delta=\partial_{xx}+\partial_{yy}$ the Laplacian operator. The periodic unsteady Stokes system reads $\partial_{t}\mathbf{u}\;-\;\nu\,\Delta\mathbf{u}\;+\;\nabla p=\mathbf{0}$ where $\nabla\!\cdot\!\mathbf{u}=$. We synthesize $200$ samples by drawing an 8-mode Fourier stream-function $\psi$, setting $\mathbf{u}_{0}=\nabla^{\!\perp}\psi$, and evolving to $T=0.1$ with $\nu=0.1$ via the analytic heat kernel. Each pair is sampled on a grid $16\times 16$, channelwise min-max scaled to $[0,1]$, and split $160{:}40$ (train test).
Table 1: Mean squared error (mean $\pm$ std. dev.) of classical vs. copresheaf architectures for learning various physics simulations.
@@
[…]
@@
For heat and advection, we leverage two transformer layers (positional encoding, 4 heads, stalk-dim $=16$) $\rightarrow$ followed by a mean pooling and linear head yielding 64D token embeddings. We train our networks using AdamW ($10^{-3}$) with a cosine LR scheduling, and batch size of two. We use 50 epochs for Heat dataset and 80 for Advection and report the results over three seeds. For Unsteady stokes, we test a compact Conv–Transformer U-Net: $2\!\rightarrow\!32$ conv encoder $\rightarrow$ two Transformer layers (4 heads, $d=32$, stalk-dim $=8$) $\rightarrow$ conv decoder back to two channels. We train it for 300 epochs using AdamW with a batch size of four. The classical baselines always use dot-product attention whereas the copresheaf variants employ learned outer-product maps.
@@ Results
As shown in Tab. 1, in heat and advection tests copresheaf attention significantly outperforms classical dot-product attention, reducing the test MSE by over $50\%$ and achieving more stable results across seed. For unsteady stokes, the copresheaf attention lowers the test MSE by $\approx 15\,\%$ and reduces variance by an order of magnitude, confirming that pair-specific linear transports capture viscous diffusion of vorticity more faithfully than standard self-attention under identical compute budgets.
@@ 5.2 Graph Classification
We evaluate whether incorporating copresheaf structure into graph neural networks (GNNs) improves performance on graph classification tasks.
@@
MUTAG dataset, a nitroaromatic compound classification benchmark consisting of 188 molecular graphs, where nodes represent atoms and edges represent chemical bonds. Each node is associated with a 7-dimensional feature vector encoding atom type, and each graph is labeled as mutagenic or non-mutagenic (2 classes). The dataset is split into 80% training and 20% testing, with experiments repeated over multiple runs to ensure robustness.
@@
We compare standard GNN models (GCN, GraphSAGE, GIN) against their copresheaf-counterparts (CopresheafGCN, CopresheafSage, CopresheafGIN) derived below. All models are two-layer networks with a hidden dimension of 32 for GCN and GraphSAGE, and 16 for GIN, followed by global mean pooling and a linear classifier to predict graph labels. The standard models (GCN (Kipf and Welling, 2017b), GraphSAGE (Hamilton et al., 2017), GIN (Xu et al., 2019)) use conventional GNN convolutions: GCN with symmetric normalization, GraphSAGE with mean aggregation, and GIN with sum aggregation. The copresheaf-enhanced models (CopresheafGCN, CopresheafSage, CopresheafGIN) augment these with learned per-edge copresheaf maps, introducing local consistency constraints via transformations. All models are trained using Adam (lr=0.01) with a batch size of 16. GCN and GIN models are trained for 100 epochs, while GraphSAGE for 50. The negative log-likelihood loss is minimized, and performance is evaluated via test accuracy. For GCN and GraphSAGE, we use 5 runs, while GIN uses 10 runs.
@@
The copresheaf structure enhances each GNN by learning a transport map $\rho_{ij}=I+\Delta_{ij}$ for each edge $(i,j)$, where $\Delta_{ij}$ is a learned transformation and $I$ is the identity matrix. The process for each model is as follows:
@@
  • CopresheafGCN. For node features $\mathbf{h}_{i},\mathbf{h}_{j}$, compute $\Delta_{ij}=\tanh(\text{Linear}([\mathbf{h}_{i},\mathbf{h}_{j}]))$, a diagonal $D\times D$ matrix. Form $\rho_{ij}=I+\Delta_{ij}$. Aggregate neighbor features as $\mathbf{h}_{i}^{\prime}=\sum_{j\in\mathcal{N}(i)}(\rho_{ij}\mathbf{h}_{j})/{\sqrt{d_{i}d_{j}}}$, where $d_{i},d_{j}$ are degrees. Combine with the self-feature: $\mathbf{h}_{i}^{\prime\prime}=(1+\epsilon)\mathbf{h}_{i}+\mathbf{h}_{i}^{\prime}$.
@@
  • CopresheafGIN. Compute $\Delta_{ij}=\tanh(\text{Linear}([\mathbf{h}_{i},\mathbf{h}_{j}]))$, a full $D\times D$ matrix, and form $\rho_{ij}=I+\Delta_{ij}$. Aggregate: $\mathbf{h}_{i}^{\prime}=\sum_{j\in\mathcal{N}(i)}(\rho_{ij}\mathbf{h}_{j})/d_{i}$. Combine: $\mathbf{h}_{i}^{\prime\prime}=(1+\epsilon)\mathbf{h}_{i}+\mathbf{h}_{i}^{\prime}$. The map $\rho_{ij}$ enables rich structural transformations.
@@ Results
Copresheaf-enhanced GNNs consistently outperform their standard counterparts on the MUTAG dataset across all three architectures (GCN, GraphSAGE, GIN), with CopresheafSage showing the highest average accu- racy (0.732) and the largest relative improvement over GraphSAGE (0.689). The learned per-edge transport maps in copresheaf models enhance the ability to capture complex graph structures, improving classification performance by enforcing local consistency. These results highlight the potential of copresheaf structures in improving GNN performance for molecular graph classification tasks.
@@ 5.3 Combinatorial Complex Classification
Finally, we assess whether incorporating copresheaf structure into transformer-based attention mechanisms improves performance on classifying higher-order, general data structures, such as combinatorial complexes (CCs). Specifically, we compare a standard transformer model (Classical) against two CT variants (CT-FC and CT-SharedLoc) on a synthetic dataset of CCs.
@@
Our synthetic dataset comprises 200 training and 50 test CCs derived from Erdős-Rényi graphs, each with 10 nodes and a base edge probability of 0.5. Triangles are added to form higher-order structures with probability $q=0.1$ for class 0 (low density) or $q=0.5$ for class 1 (high density). Nodes have 2-dimensional feature vectors, consisting of their degree and the number of triangles it participates in, with added Gaussian noise ($\mathcal{N}(0,0.1)$).
@@
All models are transformer-based classifiers with a single block, using 2 attention heads, an embedding dimension of 8, and a head dimension of 4. The model embeds 2D node features, applies attention, performs global average pooling, and uses a linear classifier to predict one of two classes (low or high triangle density). The classical model employs standard multi-head attention. The CT-FC and CT-SharedLoc models augment attention with learned per-edge transport maps to enforce local consistency. Models are trained for 4 epochs using Adam (learning rate $10^{-3}$) with a batch size of 8, minimizing cross-entropy loss. Performance is evaluated via test accuracy. Experiments are over 4 runs with different random seeds to ensure robustness.
@@ Results
Table 3: Mean and std test accuracy for CC classification.
@@
task, with CT-SharedLoc achieving the highest average accuracy (0.970) and competitive stability (std 0.010). The learned per-edge transport maps $\rho_{ij}$ enhance the model’s ability to capture higher-order structural patterns, such as triangle density, by aligning node features effectively. CT-SharedLoc’s combination of shared transport maps and local modulation yields the best performance, underscoring the value of copresheaf structures in transformer-based models for combinatorial complex classification.

### v3

@@ 5 Architectures Derived from the Copresheaf Framework
Having established the abstract copresheaf-based framework on a CC $\mathcal{X}$, we now present several concrete instantiations. Copresheaf transformers (CTs) extend the standard attention mechanism by dynamically learning linear maps $\rho_{y\to x}:\mathcal{F}(y)\to\mathcal{F}(x)$ encoding directional, anisotropic relationships between tokens, i.e., cells, in $\mathcal{X}$. Integrating these maps into attention enables CTs to capture rich, structured interactions. Copresheaf graph neural networks (CGNNs) generalize message-passing GNNs by incorporating copresheaf linear maps to model relational structures. Copresheaf convolutional networks define convolution-like operations on CCs, modeled as a Euclidean grid, using these linear maps. We present the CT construction next and leave the exact formulations of copresheaf networks, CGNNs, and copresheaf convolution layers to the appendices EF and G.
@@
Having established the abstract framework of CTNNs on a CC $\mathcal{X}$, we now introduce a concrete instantiation: the copresheaf transformer (CT) layer. This layer extends the standard attention mechanism by dynamically learning linear maps $\rho_{y\to x}:\mathbb{R}^{d_{y}}\to\mathbb{R}^{d_{x}}$ that encode both the combinatorial structure and directional, anisotropic relationships within the complex. By integrating these maps into the attention computation, the CT layer captures rich, structured interactions across $\mathcal{X}$. At layer $\ell$, each cell $x\in\mathcal{X}$ is associated with a feature $\mathbf{h}_{x}^{(\ell)}\in\mathbb{R}^{d_{x}}$, where $d_{x}$ denotes the feature dimension of cell $x$.
Definition 11 (Copresheaf self-attention).
For a fixed rank $k$ and neighborhood $\mathcal{N}_{k}$ (e.g., adjacency between $k$-cells), let $W_{q},W_{k}\in\mathbb{R}^{p\times d},W_{v}\in\mathbb{R}^{d\times d}$ denote learnable projection matrices, where $p$ is the dimension of the query and key spaces, and $d$ is the feature dimension (assumed uniform across cells for simplicity). For each $k$-cell $x\in\mathcal{X}^{k}$, copresheaf self-attention defines the message aggregation and feature update as $\mathbf{h}_{x}^{(\ell+1)}=\beta\bigl(\mathbf{h}_{x}^{(\ell)},m_{x}\bigr)$, where $m_{x}=\sum_{y\in\mathcal{N}_{k}(x)}a_{xy}\rho_{y\to x}(v_{y})$ and
@@
where $q_{x}=W_{q}h_{x}^{(\ell)},k_{x}=W_{k}h_{x}^{(\ell)},~\mathrm{and}~v_{x}=W_{v}h_{x}^{(\ell)}$. Here, the softmax normalizes over all neighbors $y^{\prime}\in\mathcal{N}_{k}(x)$ and $\rho_{y\to x}:\mathcal{F}(y)\to\mathcal{F}(x)$ is the learned copresheaf map. The update function $\beta$ is chosen to be a neural network.
Similarly, we define copresheaf cross-attention among $s$ and $t$ rank cells in $\mathcal{X}$, as well as a general algorithm for a corpresheaf transformer layer (Appendix F).
@@ 6 Experimental Evaluation
We conduct experiments on synthetic and real data in numerous settings to support the generality of our framework. These include learning physical dynamics, graph classification in homophillic and heterophilic cases and classifying higher-order complexes.
@@ 6.1 Evaluations on Physics Datasets
To verify the validity of our networks in toy setups of different phenomena, we generate a series of synthetic datasets. These include:
@@
    Advection. Similar to heat, we generate 600 realisations of $u_{t}+c\,u_{x}=0$ with $c=1$; the solution is a pure phase shift of $u_{0}$, sampled on $N=130$ points. Each pair is normalized to the interval $[0,1]$ and the dataset is split into 500:100500{:}100 train/test samples.
@@
    Unsteady stokes. Let $\mathbf{u}(x,y,t)=\bigl(u(x,y,t),\,v(x,y,t)\bigr)\in\mathbb{R}^{2}$ denote the incompressible velocity field, $p(x,y,t)\in\mathbb{R}$ the kinematic pressure, and $\nu>0$ the kinematic viscosity. Throughout, $\partial_{t}$ is the time derivative, $\nabla=(\partial_{x},\partial_{y})$ the spatial gradient, and $\Delta=\partial_{xx}+\partial_{yy}$ the Laplacian operator. The periodic unsteady Stokes system reads $\partial_{t}\mathbf{u}-\nu\Delta\mathbf{u}+\nabla p=\mathbf{0}$, where $\nabla\cdot\mathbf{u}=0$. We synthesize $200$ samples by drawing an 8-mode Fourier stream-function $\psi$, setting $\mathbf{u}_{0}=\nabla^{\!\perp}\psi$, and evolving to $T=0.1$ with $\nu=0.1$ via the analytic heat kernel. Each pair is sampled on a $16\times 16$ grid, with each channel normalized to the interval $[0,1]$, and the dataset is split into 160:40160{:}40 train/test samples.
Table 2: Mean squared error (mean $\pm$ standard deviation) of classical vs copresheaf architectures for learning various physics simulations.
@@
Classical $2.64\times 10^{-4}\;\pm\,3.50\times 10^{-5}$ $3.52\times 10^{-4}\,\pm\,7.70\times 10^{-5}$ $1.75\times 10^{-2}\,\pm\,1.32\times 10^{-3}$
Copresheaf $\mathbf{9.00\times 10^{-5}\;\pm\;7.00\times 10^{-6}}$ $\mathbf{1.20\times 10^{-4}\,\pm\,1.20\times 10^{-5}}$ $\mathbf{1.48\times 10^{-2}\,\pm\,1.48\times 10^{-4}}$
@@
For heat and advection, we use two transformer layers (positional encoding, four heads, stalk dimension equal to $16$), followed by a mean pooling and linear head yielding 64D token embeddings. We train our networks using AdamW wiht a learning rate of $10^{-3}$, cosine LR scheduling, and batch size of two. We use $50$ epochs for the heat dataset and $80$ for the advection dataset, and report the results over three seeds. For the unsteady Stokes data, we test a compact convolution-transformer U-Net consisting of a convolutional encoder with two input and $32$ output channels, followed by two transformer layers (four heads, hidden dimension equal to $32$, and stalk dimension $8$), and a convolutional decoder mapping back to two output channels. We train it for $300$ epochs using AdamW with a batch size of four. The classical baselines use dot-product attention, whereas the copresheaf variants employ learned outer-product maps.
@@ Results
As shown in Table 2, in the heat and advection tests, copresheaf attention significantly outperforms classical dot-product attention, reducing the test MSE by over $50\%$ and achieving more stable results across seeds. For unsteady Stokes, the copresheaf attention lowers the test MSE by $\approx 15\%$ and reduces variance by an order of magnitude, confirming that pair-specific linear transports capture viscous diffusion of vorticity more faithfully than standard self-attention under identical compute budgets.
@@ 6.2 Graph Classification
We evaluate whether incorporating copresheaf structure into GNNs improves performance on graph classification tasks.
@@
MUTAG dataset, a nitroaromatic compound classification benchmark consisting of $188$ molecular graphs, where nodes represent atoms and edges represent chemical bonds. Each node is associated with a $7$-dimensional feature vector encoding atom type, and each graph is labeled as mutagenic or non-mutagenic (two classes). The dataset is split into $80\%$ train and $20\%$ test samples.
@@
We compare standard GNN models (GCN, GraphSAGE, GIN) against their copresheaf-counterparts (CopresheafGCN, CopresheafSage, CopresheafGIN) derived below. All models are two-layer networks with a hidden dimension of $32$ for GCN and GraphSAGE, and $16$ for GIN, followed by global mean pooling and a linear classifier to predict graph labels. The standard models (GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GIN (Xu et al., 2019)) use conventional GNN convolutions: GCN with symmetric normalization, GraphSAGE with mean aggregation, and GIN with sum aggregation. The copresheaf-enhanced models augment these with learned per-edge copresheaf maps, introducing local consistency constraints via transformations. All models are trained using Adam with a learning rate of $0.01$, and a batch size of $16$. GCN and GIN models are trained for $100$ epochs, while GraphSAGE for $50$. The negative log-likelihood loss is minimized, and performance is evaluated via test accuracy. For GCN and GraphSAGE, we use five runs, while GIN uses ten runs.
@@
The copresheaf structure enhances each GNN by learning a transport map $\rho_{ij}=I+\Delta_{ij}$ for each edge $(i,j)$, where $\Delta_{ij}$ is a learned transformation and $I$ is the identity matrix. In what follows $D$ represents the dimension of the input feature in where we apply the copresheaf maps $\rho_{ij}$. Denote by $\mathbf{h}_{i};\mathbf{h}_{j}\,]$ to the concatenation of the two node feature vectors $\mathbf{h}_{i}$ and $\mathbf{h}_{j}$ along their feature dimension. The process for each model is as follows:
@@
  • CopresheafGCN. For node features $\mathbf{h}_{i},\mathbf{h}_{j}$, compute $\Delta_{ij}=\tanh(\mathrm{Linear}([\mathbf{h}_{i};\mathbf{h}_{j}]))$, take its diagonal to get a diagonal $D\times D$ matrix with $D=7$. Form $\rho_{ij}=I+\Delta_{ij}$. Aggregate neighbor features as $\mathbf{h}_{i}^{\prime}=\sum_{j\in\mathcal{N}(i)}\frac{1}{\sqrt{d_{i}d_{j}}}\;\rho_{ij}\mathbf{h}_{j},$ where $d_{i},d_{j}$ are the degrees of nodes $i$ and $j$, respectively. Combine with the self-feature: $\mathbf{h}_{i}^{\prime\prime}=(1+\epsilon)\mathbf{h}_{i}+\mathbf{h}_{i}^{\prime}.$
@@
  • CopresheafGIN. Compute $\Delta_{ij}=\tanh(\text{Linear}([\mathbf{h}_{i},\mathbf{h}_{j}]))$, a full $D\times D$ matrix, and form $\rho_{ij}=I+\Delta_{ij}$. Aggregate: $\mathbf{h}_{i}^{\prime}=\sum_{j\in\mathcal{N}(i)}(\rho_{ij}\mathbf{h}_{j})/d_{i}$. Combine: $\mathbf{h}_{i}^{\prime\prime}=(1+\epsilon)\mathbf{h}_{i}+\mathbf{h}_{i}^{\prime}$.
@@ Results
On MUTAG, copresheaf-enhanced GNNs consistently outperform their standard versions across GCN, GraphSAGE, and GIN. CopresheafSAGE achieves the highest average accuracy ($0.732$) and the largest relative gain over the GraphSAGE baseline ($0.689$). Learned per-edge transport maps better capture complex structure and enforce local consistency, improving classification. These results demonstrate the promise of copresheaf structures for molecular graph classification.
@@ 6.3 Combinatorial Complex Classification
Finally, we assess whether incorporating copresheaf structure into transformer-based attention mechanisms improves performance on classifying higher-order, general data structures, such as CCs. Specifically, we compare a classical transformer model against two CT variants (CT-FC and CT-SharedLoc) on a synthetic dataset of CCs.
@@
Our synthetic dataset comprises $200$ training and $50$ test CCs derived from Erdős-Rényi graphs, each with $10$ nodes and a base edge probability of $0.5$. Triangles are added to form higher-order structures with probability $q=0.1$ for class 0 (low density) or $q=0.5$ for class $1$ (high density). Each node has 2D feature vectors, consisting of its degree and the number of triangles it participates in, with added Gaussian noise $\mathcal{N}(0,0.1)$.
@@
All models are transformer-based classifiers with a single block, using two attention heads, an embedding dimension of $8$, and a head dimension of $4$. The model embeds 2D node features, applies attention, performs global average pooling, and uses a linear classifier to predict one of two classes (low or high triangle density). The classical model employs standard multi-head attention. The CT-FC and CT-SharedLoc models augment attention with learned per-edge transport maps to enforce local consistency. Models are trained for four epochs using Adam with a learning rate of $10^{-3}$ and a batch size of $8$, minimizing cross-entropy loss. Performance is evaluated via test accuracy. Experiments are over four runs with different random seeds to ensure robustness.
@@ Results
Table 4: Mean $\pm$ std test accuracy for CC classification.
@@
task, with CT-SharedLoc achieving the highest average accuracy ($0.970$) and competitive stability (std $0.010$). The learned per-edge transport maps $\rho_{ij}$ enhance the model’s ability to capture higher-order structural patterns, such as triangle density, by aligning node features effectively. CT-SharedLoc’s combination of shared transport maps and local modulation yields the best performance, showcasing the value of copresheaf structures in transformer-based models for CC classification.